# PLAME: Lightweight MSA Design Advances Protein Folding From Evolutionary Embeddings

### v2

@@ 1 Introduction
Understanding complex and variable protein structures is central to target identification, validation, and drug–target interaction studies in drug design baker2001protein ; khoury2014protein ; jumper2021alphafold . Recent advances such as AlphaFold have transformed structural biology, achieving near-experimental accuracy across a wide range of proteins and complexes jumper2021alphafold ; ahdritz2024openfold ; abramson2024accurate . However, most state-of-the-art folding pipelines depend critically on evolutionary information encoded in Multiple Sequence Alignments (MSAs) jumper2021alphafold ; abramson2024accurate . As a consequence, their accuracy tracks the quality and depth of the available MSAs, leading to failure modes for low-homology families and orphan proteins where evolutionary neighbors are sparse or absent kwon2021assessment ; webb2016comparative . In practice, while MSA depth (e.g., effective sequence count) often correlates with predicted confidence (such as pLDDT), this relationship becomes highly unstable in sparse regimes where a few noisy or misaligned sequences can dominate the signal.
@@
Historically, two classes of techniques have been used to mitigate weak homology. Physics-based modeling seeks low-energy conformations under hand-crafted or learned force fields, but is often computationally intensive and limited by approximations in the energy landscape rohl2004rosetta ; cornell1995amber . Template-based approaches leverage homology detection and profile–profile alignment to transfer structural priors from known folds hildebrand2009fast ; finn2011hmmer , yet degrade in regimes with poor evolutionary signal and are thus ill-suited to orphan proteins. These limitations have motivated data-driven strategies focused on improving the MSA itself rather than only the downstream folding network.
@@
Recent MSA design methods fall broadly into two lines (Figure 1). Sequence-space inpainting approaches (e.g., MSA Generator, EvoGen) complete or augment partial alignments by learning patterns directly in the discrete sequence space, aiming to reconstruct evolutionary constraints from observed MSAs zhang2023enhancing ; zhang2022few . Prompt-based conditional generation (e.g., MSAGPT, EvoDiff) uses pretrained models conditioned on MSA-style prompts to synthesize additional sequences chen2024msagpt ; alamdari2023protein . These approaches can deepen alignments and improve folding when homologs exist. An orthogonal line of work bypasses explicit MSAs by constructing implicit evolutionary representations from single sequences via large protein language models, as in ESMFold lin2023evolutionary . While MSA-free models avoid the homology bottleneck, they also forgo explicit template usage and enhanced homology signals, which can cap ultimate folding accuracy in challenging regimes.
Figure 1: Taxonomy of MSA designers. Prior work models the discrete MSA distribution via sequence inpainting or prompt-based generation. PLAME operates in an evolutionary embedding space to de novo generate MSAs without requiring sequence prompts.
Despite progress, two gaps remain for MSA design aimed at improving folding. (i) Biased supervision: Methods trained primarily on extant MSA databases inherit coverage biases toward well-studied families, offering limited guidance for low-homology and orphan proteins. This restricts the learned evolutionary manifold and hampers generalization exactly where better MSAs are most needed. (ii) Weak alignment–folding linkage: There is limited understanding—and few lightweight metrics—connecting MSA characteristics to downstream folding outcomes. As a result, generation objectives (e.g., likelihood in sequence space) can be misaligned with what most improves structure accuracy. While fine-tuning folding models with generated data (e.g., MSAGPT) partially closes the loop chen2024msagpt , the compute cost is high and does not yield general, model-agnostic criteria for selecting helpful MSAs. Closing this loop by retraining or adapting a folding model for each candidate MSA is prohibitive at scale, preventing iterative design-and-select cycles across large target sets.
We introduce PLAME (PLm Aligner for MSA Enhancement), a lightweight MSA designer targeted at low-homology and orphan proteins. PLAME departs from discrete sequence modeling and instead generates in an evolutionary embedding space derived from pretrained protein language models, then decodes to MSAs for downstream folding. This design enables PLAME to synthesize plausible evolutionary neighborhoods even when observed homologs are scarce, while remaining compatible with template-using pipelines.
  • Enriched evolutionary spaces. PLAME conditions on evolutionary embeddings from pretrained protein language models wang2024maape ; mcwhite2023leveraging ; hong2024fast , providing richer contextual signals than sparse MSAs for low-homology targets. We introduce a conservation–diversity loss that explicitly balances agreement at conserved positions with appropriate sequence variability, steering generation toward biologically plausible and alignment-friendly MSAs.
  • Linking alignment to folding. We propose a complementary sequence-quality assessment metric that captures alignment properties predictive of folding gains, and we operationalize these insights with HiFiAD (High-Fidelity Appropriate Diversity), an MSA selection strategy that filters candidates to maximize downstream structure accuracy.
  • Broad, model-agnostic validation. Across challenging low-homology and orphan datasets, PLAME improves folding accuracy with both AlphaFold2 and AlphaFold3. Ablations isolate the effects of the conservation–diversity loss and HiFiAD on MSA quality and folding outcomes, and case studies spanning general and de novo proteins elucidate how specific MSA characteristics modulate AlphaFold confidence and error modes.
@@ 2 Related Works
Protein structure prediction methods fall into three main categories: physics-based, homology-based, and deep learning approaches. Physics-based methods, such as AMBER and CHARMM, use molecular physics and energy optimization to simulate protein folding cornell1995amber ; brooks2009charmm . While offering detailed folding insights, they are computationally expensive and sensitive to initial conditions, often yielding suboptimal results karplus2002molecular ; freddolino2010challenges ; pande2010everything . Homology modeling tools, like Rosetta and HHpred, use MSAs and evolutionary data to predict structures by refining templates from known experimental structures rohl2004rosetta ; hildebrand2009fast . These methods perform well with suitable templates but struggle with orphan proteins and low-homology families webb2016comparative ; baker2001protein . Deep learning-based methods, such as AlphaFold2 and OmegaFold, use advanced neural architectures and protein templates to achieve near-experimental accuracy with greater speed and scalability jumper2021alphafold ; abramson2024accurate ; wu2022high . Despite their success, they still depend on high-quality MSAs and struggle with low-homology proteins.
@@
Building on AlphaFold’s success, researchers have developed methods to refine specific modules, aiming to improve accuracy or efficiency. These advancements can be grouped into three main categories. The first category focuses on homology expansion techniques, such as MMSeq2 and DeepMSA2, which expand the evolutionary search space to enhance prediction accuracy. However, these methods often slow down inference despite their modest performance gains johnson2010hidden ; steinegger2017mmseqs2 ; zheng2024improving ; lee2024petabase . The second category targets search acceleration, with methods like ColabFold and ESMFold bypassing the MSA search process to enhance computational efficiency. However, this speedup often results in incomplete evolutionary data, potentially reducing prediction accuracy lin2023evolutionary ; mirdita2022colabfold . The third category leverages generative models to capture protein homology and augment input data, especially for orphan proteins and low-homology families. While promising in specific scenarios, these models struggle with extremely limited evolutionary signals, and their artificial sequences often deviate from traditional MSA distributions, limiting broader applicability alamdari2023protein ; zhang2022few ; zhang2023enhancing ; chen2024msagpt .
@@ 3.1 Problem Formulation
where the augmented MSAs are composed of original MSAs $\mathbf{M}$ and generated MSAs $\mathbf{M}^{\prime}$, denoted as $\mathbf{M}_{\text{aug}}=\{\mathbf{M},\mathbf{M}^{\prime}\}$. The quality of the enhanced structures is evaluated using several metrics, including RMSD, TM-score, and pLDDT.
@@ 3.2 Model Architecture
PLAME employs an encoder-decoder transformer architecture similar to MSA Transformer rao2021msa , with adjustments to the T5 block structure vaswani2017attention . The encoder and decoder incorporate additional row-wise and column-wise attention mechanisms, designed to better capture evolutionary patterns in MSA data (detailed in Fig 2). Additional mechanisms are introduced as follows.
@@
The row attention mechanism models inter-sequence dependencies within the evolutionary space, summarizing sequence relationships across the depth of the input representation. Given an input tensor $\mathbf{H}_{\text{enc}}\in\mathbb{R}^{L\times N\times D\times h}$, where $L$ is the sequence length, $N$ is the number of MSAs, $D$ is the depth, and $h$ is the hidden dimension, a global representation $\mathbf{H}r$ is computed by averaging the hidden states along the depth dimension:
$\mathbf{H}_{r}=\frac{1}{D}\sum_{d=1}^{D}\mathbf{H}_{\text{enc}}^{d},\quad\mathbf{H}_{r}\in\mathbb{R}^{L\times N\times h}.$ (4)
Here, $\mathbf{H}_{r}$ encodes the evolutionary space, facilitating cross-attention during decoding. The cross-row attention (Row-Attn) is defined as:
@@
where $\mathbf{Q}_{r},\mathbf{K}_{r},\mathbf{V}_{r}$ denote the Query, Key, and Value matrices, respectively.
@@
Column attention captures conservation patterns across columns in the MSA, focusing on evolutionary signals at specific positions. To implement this mechanism, the input representation matrix $\mathbf{X}_{\text{dec}}\in\mathbb{R}^{D\times N\times L\times h}$ is first transposed to $\mathbf{X}_{\text{dec}}^{\top}\in\mathbb{R}^{L\times N\times D\times h}$. The self and cross-column attention mechanisms operate on this transformed representation.
For cross-column attention, the query, key, and value matrices $\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}}$ and their corresponding projection weights $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are defined as:
$\mathbf{Q}_{\text{c}}=\mathbf{X}_{\text{dec}}^{\top}\mathbf{W}_{q},\quad\mathbf{K}_{\text{c}}=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{k},\quad\mathbf{V}_{\text{c}}=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{v}.$ (6)
The column attention is computed as:
$\text{Col-Att}(\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}})=\left(\text{softmax}\left(\frac{\mathbf{Q}_{\text{c}}\mathbf{K}_{\text{c}}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{\text{c}}\right)^{\top}$ (7)
@@
In our implementation, ESM-2 ($\mathbf{g}_{\phi}$) encodes the query sequence $\mathbf{s}$ into high-dimensional evolutionary embeddings, denoted as $\mathbf{H}_{\text{input}}$. These embeddings are processed by the encoder through $N$ layers of modified T5 blocks, iteratively refining contextualized representations:
@@
The final encoder output, $\mathbf{H}^{(N)}$, captures comprehensive sequence context and is passed to the decoder for autoregressive generation. The decoder generates tokens sequentially, conditioned on the encoder output $\mathbf{H}_{\text{Enc}}^{(N)}$ and previously generated tokens:
@@
The T5 blocks with row and column attention enable the decoder to compute output embeddings, which are then passed through a softmax layer to produce token probabilities for the next position.
@@ 3.3 Conservation-Diversity Training Loss
We introduce a position-aware causal inference approach for diverse MSA generation, integrating a PSSM-Weighted Cross-Entropy Loss and a Diversity Regularization Loss to balance focus on conserved regions with sampling diversity.
PSSM-Weighted Cross Entropy (PCE) Loss
The PSSM-Weighted Cross Entropy Loss emphasizes accurate predictions in conserved regions of the MSA, which are critical for maintaining protein structure and function. For a single sequence, it is defined as:
@@
where $L$ is the sequence length, $y_{l}$ is the true label at position $l$, and $p(y_{l}\mid y_{<l})$ is the predicted probability of $y_{l}$.
The position-specific weights $w_{l}$ are derived from the Position-Specific Scoring Matrix (PSSM) and reflect the conservation level at each position. These weights are normalized to the range $1-\delta,1+\delta]$, where $\delta$ controls sensitivity to conservation. Specifically,
@@
Diversity Regularization (DIRE) Loss
The Diversity Regularization Loss promotes biological diversity in MSAs by maximizing the entropy of predicted amino acid distributions. For a single sequence, the entropy at position $l$ is calculated as:
$H_{l}=-\sum_{a\in\mathcal{A}}p(a\mid y_{<l})\log p(a\mid y_{<l}),$ (13)
where $p(a\mid y_{<l})$ is the predicted probability of amino acid $a$ at position $l$, and $\mathcal{A}$ is the set of all amino acids. To encourage diversity across a batch of $N$ sequences, we compute the average entropy over all positions and sequences:
$\mathcal{L}_{\text{diversity}}=-\frac{1}{N}\sum_{j=1}^{N}\frac{1}{L_{j}}\sum_{l=1}^{L_{j}}H_{l}^{(j)},$ (14)
where $L_{j}$ is the length of sequence $j$, and $H_{l}^{(j)}$ is the entropy at position $l$ in sequence $j$. This loss encourages the model to capture the natural amino acid diversity in homologous sequences.
@@
where $\alpha\in[0,1]$ controls the trade-off between the two terms. We set $\alpha=0.9$ in our experiments, prioritizing conservation while maintaining sufficient diversity. This design ensures the generated MSAs accurately capture conserved regions and retain natural sequence variability. Our theoretical analysis in Appendix A demonstrates that PCE Loss enhances the model’s understanding of evolutionary information (MSA profile), while DIRE Loss functions as a regularizer to prevent neglect of variable regions.

### v3

# Lightweight MSA Design Advances Protein Folding From Evolutionary Embeddings

@@ 1 Introduction
Understanding complex and dynamic protein structures is fundamental to target identification, validation, and drug-target interaction studies in drug design (baker2001protein, ; khoury2014protein, ). Recent advances such as AlphaFold have revolutionized structural biology, achieving near-experimental accuracy across a broad spectrum of proteins and complexes (jumper2021alphafold, ; ahdritz2024openfold, ; abramson2024accurate, ). However, most state-of-the-art folding pipelines heavily rely on evolutionary information encoded within multiple sequence alignments (MSAs) (lin2023evolutionary, ; abramson2024accurate, ). Consequently, their accuracy is highly correlated with the quality and depth of available MSAs. This dependency creates failure modes in low-homology families and orphan proteins (those lacking or having few evolutionary neighbors) (kwon2021assessment, ; webb2016comparative, ), where even small amounts of noisy or misaligned sequences can dominate the signal.
@@
Historically, two primary classes of techniques have been developed to address weak homology. Physics-based modeling searches for low-energy conformations in energy space through handcrafted or learned force fields, but is often computationally intensive and limited by approximations in the energy landscape (rohl2004rosetta, ; cornell1995amber, ). Template-based methods leverage homology detection and profile-profile alignment to transfer structural priors from known folds to novel sequences (hildebrand2009fast, ; finn2011hmmer, ), but suffer degraded performance in the absence of evolutionary signals, making them unsuitable for orphan proteins. These limitations have motivated a shift toward data-driven strategies that focus on improving the MSA itself rather than solely the downstream folding networks.
@@
Recent MSA design approaches can be broadly categorized into two paradigms (Figure 1). Sequence-space inpainting methods (e.g., MSA Generator, EvoGen) directly learn patterns in discrete sequence space to augment partial alignments, aiming to reconstruct evolutionary constraints from existing MSAs (zhang2023enhancing, ; zhang2022few, ). Prompt-based conditional generation approaches (e.g., MSAGPT, EvoDiff) utilize pre-trained models to synthesize additional sequences under MSA-style prompts (chen2024msagpt, ; alamdari2023protein, ). These methods can deepen alignments and improve folding accuracy when homologous sequences exist. An orthogonal line of research bypasses explicit MSA construction by building implicit evolutionary representations from single sequences through large protein language models (PLMs), as demonstrated by ESMFold (lin2023evolutionary, ). While MSA-free models avoid the homology bottleneck, they also forgo explicit template usage and enhanced homology signals, which may limit ultimate folding accuracy in challenging scenarios.
Despite existing progress, two critical gaps remain in structure prediction for low-homology proteins. (i) Supervision bias: Methods trained on existing MSA databases inherit biases toward well-studied families, limiting effectiveness for low-homology and orphan proteins. (ii) Weak alignment-folding correlation: Current approaches lack lightweight metrics linking MSA characteristics to folding outcomes. Sequence-based generative objectives may not align with factors that improve structural accuracy, while existing solutions like fine-tuning folding models (chen2024msagpt, ) are computationally expensive and lack universal applicability.
In this study, we propose PLAME, motivated by the critical need to enhance structure prediction for low-homology proteins where traditional MSA-based approaches fail due to insufficient evolutionary signals. Our approach makes the following key contributions:
  1. 1.
    Embedding-space MSA generation with conservation-diversity optimization: Inspired by PLMs’ success in MSA-related tasks (hong2024fast, ; wang2024maape, ; mcwhite2023leveraging, ), we develop the first MSA designer that generates auto-regressively within the evolutionary embedding space of pre-trained PLMs rather than discrete sequences (Fig2). We further propose a novel conservation-diversity loss that captures conserved regions while extracting diverse variants from ESM embeddings with theoretical guarantee (AppendixA). The lightweight design enables PLAME to synthesize evolutionary neighborhoods even with scarce homologous sequences, achieving up to three orders of magnitude speedup while maintaining template compatibility (Table 4).
  2. 2.
    HiFiAD: A principled MSA quality assessment framework: To address the current weak alignment-folding correlation problem, we propose High-Fidelity Appropriate Diversity (HiFiAD), a lightweight algorithm for MSA filtering that simultaneously considers site-wise conservation and inter-MSA diversity. This provides the first model-agnostic, computationally efficient criterion for selecting high-quality alignments that directly correlate with improved folding outcomes.
  3. 3.
    Comprehensive validation across challenging scenarios: On challenging low-homology and orphan datasets, PLAME consistently improves folding accuracy in both AlphaFold2 and AlphaFold3, performing similarly to DHR (hong2024fast, ), AI-based MSA searching approach. In ablation studies, HiFiAD demonstrates performance gains across all baselines (Table4). Moreover, case studies on general and de novo proteins further demonstrate PLAME’s generalizability while providing novel perspectives on structure enhancement from an MSA design standpoint (Table8). PLAME offers new insights and possibilities for folding enhancement through principled MSA optimization.
@@ 2 Related Works
Protein structure prediction methods fall into three main categories: physics-based, homology-based, and deep learning approaches. Physics-based methods, such as AMBER and CHARMM, use molecular physics and energy optimization to simulate protein folding (cornell1995amber, ; brooks2009charmm, ). While offering detailed folding insights, they are computationally expensive and sensitive to initial conditions, often yielding suboptimal results (karplus2002molecular, ; freddolino2010challenges, ; pande2010everything, ). Homology modeling tools, like Rosetta and HHpred, use MSAs and evolutionary data to predict structures by refining templates from known experimental structures (rohl2004rosetta, ; hildebrand2009fast, ). These methods perform well with suitable templates but struggle with orphan proteins and low-homology families (webb2016comparative, ; baker2001protein, ). Deep learning-based methods, such as AlphaFold2 and OmegaFold, use advanced neural architectures and protein templates to achieve near-experimental accuracy with greater speed and scalability (jumper2021alphafold, ; abramson2024accurate, ; wu2022high, ). Despite their success, they still depend on high-quality MSAs and struggle with low-homology proteins.
@@
Building on AlphaFold’s success, researchers have developed methods to refine specific modules, aiming to improve accuracy or efficiency. These advancements can be grouped into three main categories. The first category focuses on homology expansion techniques, such as MMSeq2 and DeepMSA2, which expand the evolutionary search space to enhance prediction accuracy. However, these methods often slow down inference despite their modest performance gains (johnson2010hidden, ; steinegger2017mmseqs2, ; zheng2024improving, ; lee2024petabase, ). The second category targets search acceleration, with methods like ColabFold and ESMFold bypassing the MSA search process to enhance computational efficiency. However, this speedup often results in incomplete evolutionary data, potentially reducing prediction accuracy (lin2023evolutionary, ; mirdita2022colabfold, ). The third category leverages generative models to capture protein homology and augment input data, especially for orphan proteins and low-homology families. While promising in specific scenarios, these models struggle with extremely limited evolutionary signals, and their artificial sequences often deviate from traditional MSA distributions, limiting broader applicability (alamdari2023protein, ; zhang2022few, ; zhang2023enhancing, ; chen2024msagpt, ).
@@ 3.1 Problem Formulation
where the augmented MSAs are composed of original MSAs $\mathbf{M}$ and generated MSAs $\mathbf{M}^{\prime}$, denoted as $\mathbf{M}_{\text{aug}}=\{\mathbf{M},\mathbf{M}^{\prime}\}$. The quality of the enhanced structures is evaluated using several metrics, including RMSD, TM-score, and pLDDT (See details in Section4).
@@ 3.2 Model Architecture
PLAME employs an encoder-decoder transformer architecture similar to MSA Transformer (rao2021msa, ), with adjustments to the T5 block structure (vaswani2017attention, ). The encoder and decoder incorporate additional row-wise and column-wise attention mechanisms to better capture evolutionary patterns in MSA data (detailed in Fig 2), which is similarly applied in MSAGenerator (zhang2023enhancing, ) and MSAGPT (chen2024msagpt, ). Additional mechanisms are introduced as follows.
@@
Row attention models inter-sequence dependencies by summarizing evolutionary relationships across MSA depth. Given input $\mathbf{H}_{\text{enc}}\in\mathbb{R}^{L\times N\times D\times h}$, we compute a global representation by depth-wise averaging:
$\mathbf{H}_{r}=\frac{1}{D}\sum_{d=1}^{D}\mathbf{H}_{\text{enc}}^{d}\in\mathbb{R}^{L\times N\times h},$ (4)
where $\mathbf{H}_{r}$ encodes the evolutionary space for cross-attention during decoding:
@@
Column attention captures positional conservation patterns across MSA columns. We transpose the decoder input $\mathbf{X}_{\text{dec}}\in\mathbb{R}^{D\times N\times L\times h}$ to $\mathbf{X}_{\text{dec}}^{\top}$ and compute cross-column attention with:
$\mathbf{Q}_{\text{c}}=\mathbf{X}_{\text{dec}}^{\top}\mathbf{W}_{q},\quad\mathbf{K}_{\text{c}}$ $=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{k},\quad\mathbf{V}_{\text{c}}=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{v},$ (6)
$\text{Col-Att}(\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}})$ $=\left(\text{softmax}\left(\frac{\mathbf{Q}_{\text{c}}\mathbf{K}_{\text{c}}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{\text{c}}\right)^{\top}.$ (7)
@@
ESM2 (lin2023evolutionary, ) encodes the query sequence $\mathbf{s}$ into evolutionary embeddings $\mathbf{H}_{\text{input}}$. The encoder processes these through $N$ modified T5 layers:
@@
The decoder autoregressively generates tokens conditioned on encoder output and previous tokens:
@@
Output embeddings are passed through softmax to produce token probabilities.
@@ 3.3 Conservation-Diversity Training Loss
We propose a position-aware causal inference approach for diverse MSA generation, integrating a PSSM-Weighted Cross-Entropy (PCE) Loss and a DIversity REgularization (DIRE) Loss to balance focus on conserved regions with sampling diversity.
PCE Loss
The PCE Loss emphasizes accurate predictions in conserved regions of the MSA, which are critical for maintaining protein structure and function. For a single sequence, it is defined as:
@@
where $L$ denotes sequence length, $y_{l}$ denotes the amino acid at site $l$, and $p(y_{l}\mid y_{<l})$ denotes the predicted discrete probability distribution of $y_{l}$.
The position-specific weights $w_{l}$ are derived from the Position-Specific Scoring Matrix (PSSM) (henikoff1994position, ) and reflect the conservation level at each position. These weights are normalized to the range $1-\delta,1+\delta]$, where $\delta$ controls sensitivity to conservation. Specifically,
@@
DICE Loss
The DIRE loss promotes sequence diversity by maximizing amino acid entropy:
$\mathcal{L}_{\text{DIRE}}=-\frac{1}{N}\sum_{j=1}^{N}\frac{1}{L_{j}}\sum_{l=1}^{L_{j}}H_{l}^{(j)},$ (13)
where $H_{l}^{(j)}=-\sum_{a\in\mathcal{A}}p(a\mid y_{<l})\log p(a\mid y_{<l})$ is the entropy at position $l$ in sequence $j$, and $\mathcal{A}$ is the set of all amino acids.
@@
with $\alpha=0.9$ prioritizing conservation while maintaining variability. Our theoretical analysis in AppendixA demonstrates that PCE Loss enhances the model’s understanding of MSA profile, while DIRE Loss functions as a regularizer to prevent neglect of variable regions.